OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. y = x + 2 y = -x - 2 y = -x + 4 y = x - 4

OpenStudy (anonymous):

y = -x+4 any line parallel to this should be y= -x+c where c is a constant it passes through -3,1 so putting x=-3 and y=1 we get, 1=3+c c=1-3 = -2 so the line is y=-x-2 which is the 2nd option

OpenStudy (shamim):

agree with @Koel